OpenStudy (anonymous):

STUCK FOR DAYS!!!! NEED HELP ASAP!!!!! MEDALS!!!!! FAN-ING!!!!! WHATEVER!!! I'LL DO ANYTHING FOR HELP!!!!!!!

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price A.) f(x) = 25(1.08)x B.) x f(x) 0 22.00 1 21.56 2 21.13 C.) The initial price of this stock is $30, but it has been increasing 4% each day. 1.) Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. 2.) Graph the price function for Stock D. Explain the key features of the graph with complete sentences. 3.) A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). 4.) Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

merry christmas! :D

OpenStudy (anonymous):

I need help with math not socializing

OpenStudy (anonymous):

You have to create one, that is what these questions are helping you do

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@kayla_b_stephens @ninabelkaniya

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@baby456

OpenStudy (anonymous):

@hayatsaed

OpenStudy (anonymous):

@mattdogg @luka1998

undeadknight26 (undeadknight26):

C.) The initial price of this stock is $30, but it has been increasing 4% each day. f(x) = 30(1.4)x

OpenStudy (anonymous):

Thanks @sleepyjess

undeadknight26 (undeadknight26):

1.) Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. choose a number between 25 and 30 please.

OpenStudy (anonymous):

@MERTICH said that it had to be between 22 and 25

OpenStudy (anonymous):

So 23

undeadknight26 (undeadknight26):

no it has to be between the initial start prices which were 25 and 30. now pick a number between 1.08 and 1.4

OpenStudy (anonymous):

26 and then 1.3

undeadknight26 (undeadknight26):

ok so: f(x) = 26(1.3)x that is your equation.

OpenStudy (anonymous):

That's it for number one!?!?!!?

undeadknight26 (undeadknight26):

2.) Graph the price function for Stock D. Explain the key features of the graph with complete sentences. to do this we would need to solve the equation and graph it do u know how to do this?

undeadknight26 (undeadknight26):

yes that is it.

OpenStudy (anonymous):

f(x) = 33.8x??

undeadknight26 (undeadknight26):

did you solve like this: f(1) = 26(1.3)1 f(1) = 33.8? f(2) = 26(1.3)2 f(2) = 67.6 ?

undeadknight26 (undeadknight26):

i believe so...then the x is the days and the y is the amount(or whatever :D)

undeadknight26 (undeadknight26):

do you understand so far?

undeadknight26 (undeadknight26):

3.) A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). for this you would just do this. 1 + -6 2 + -6 3 + -6 @jdoe0001 is this right?

OpenStudy (anonymous):

Yes I solved it like that

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

So can you explain to me how you got that last one??

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

This one? 4.) Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

Well yes, but the one I'm talking about is your answer for number 3

OpenStudy (anonymous):

@jdoe0001 Can you help me with number 3 and number 4??

undeadknight26 (undeadknight26):

well since g(x) = -6 and we were graphing it per day so the first day would be 1 + -6 the second day would be 2 + -6 and the third day would be 3 + -6...this is what i need jdoes help with...

OpenStudy (anonymous):

Well if you were graphing it by day then wouldn't it be g(x) = 1(-6), g(x) = 2(-6), so on and so forth??

OpenStudy (jdoe0001):

hmm what's stock D anyway?

OpenStudy (jdoe0001):

well, the function for stock D that is

undeadknight26 (undeadknight26):

no because they told me: A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). We have to use: f(x) + g(x) so f(x) + -6.

undeadknight26 (undeadknight26):

we made this: f(x) = 26(1.3)x

OpenStudy (jdoe0001):

\(\bf f(x) = 26(1.3)x\qquad g(x)=-6\\ \quad \\ f(x)-g(x)=[26(1.3)x]-6\)

OpenStudy (jdoe0001):

hmmm rather \(\bf f(x) = 26(1.3)x\qquad g(x)=-6\\ \quad \\ f(x)+g(x)=[26(1.3)x]-6\)

undeadknight26 (undeadknight26):

ohh ok... so: F(1) + g(-6) = 33.8 - 6? F(1) + g(-6) = 27.8?

OpenStudy (anonymous):

Yea, that makes sense

OpenStudy (jdoe0001):

\(\bf f(x)+g(x)=[26(1.3)x]-6\qquad f(1)+g(1)=[26(1.3)1]-6\)

OpenStudy (jdoe0001):

which will make it 33.8-6

OpenStudy (anonymous):

Ok so that's the answer?

OpenStudy (jdoe0001):

or the 1st day for the new function at stock D, will be 33.8-6

OpenStudy (jdoe0001):

well.. for 3, yes, that's how Gordon will see how it's created, by adding both functions

OpenStudy (anonymous):

What about 4

OpenStudy (jdoe0001):

well. you're supposed to also graph it

OpenStudy (jdoe0001):

once you graph the new function, 4 is just nothing but "translations" or "transformations" or "shifts" of THAT graph

OpenStudy (anonymous):

Wait what???

OpenStudy (jdoe0001):

once you have graph for f(x)+g(x), or the new function, the lines in 4 just shift it or move it about

OpenStudy (anonymous):

How would I "move" it, like I don't understand what you're saying

OpenStudy (jdoe0001):

have you covered "functions translations or transformations" yet? or function shifts

OpenStudy (anonymous):

I think so

OpenStudy (jdoe0001):